Thanks so very much! I need your help!
factor the quadratic expression completely, and find the roots of the expression.
20x^2 + 13x + 2

Answer by narayaba(40)


20x^2 + 13x + 2 = 0 of the form (ax^2 + bx + c = 0)
The product a*c = 2*20 = 40
sum = b = 13
5 times 8 is 40 5+8 = 13
hence 20x^2 + 13x + 2 can be factored as 20x^2 + 5x + 8x + 2
5x (4x + 1) + 2(4x + 1 )
(5x+2)(4x+1) = 0
x = -2/5 or x = -1/4